OpenStudy (anonymous):

can (9r-6)/(3r^2-2) be simplified?

OpenStudy (anonymous):

woops it's (9r-6)/(3r^2-2r)

OpenStudy (anonymous):

factor out an r on the bottom?

OpenStudy (anonymous):

\[\frac{ 9r-6 }{ r(3r -2)} = \frac{ 9}{ (3r-2) }-\frac{ 6 }{ 3r ^{2} -2r}\] I was thinking this but I don't think this will work.

OpenStudy (anonymous):

Oh I see it now

OpenStudy (anonymous):

easier than it seems.

OpenStudy (anonymous):

you keep going and the answer is 3/r

OpenStudy (anonymous):

\[\frac{ 3(3r-2) }{ r(3r-2) }\]

OpenStudy (anonymous):

factor out a 3 in the numerator and an r in the denominator

OpenStudy (anonymous):

haha i got it after taking the r out yup the 3r-2 cancels so you're left with 3/r

OpenStudy (anonymous):

there you go